The change in a pocket totals $1.00 and is composed of 6 coins. The number of quarters is the equal to the sum of the nickels and dimes. How many of each coin is in the pocket?
Answer by Alan3354(69443)

The change in a pocket totals $1.00 and is composed of 6 coins. The number of quarters is the equal to the sum of the nickels and dimes. How many of each coin is in the pocket?
---------------
q + n + d = 6
q = n + d
--> q + q = 6
q = 3 quarters
-------------
25q + 5n + 10d = 100
75 + 5n + 10d = 100
n + 2d = 5
--> 1 nickel
2 dimes